# IDEA: An Inverse Domain Expert Adaptation Based Active DNN IP Protection Method

### v1

@@ Abstract
Illegitimate reproduction, distribution and derivation of Deep Neural Network (DNN) models can inflict economic loss, reputation damage and even privacy infringement. Passive DNN intellectual property (IP) protection methods such as watermarking and fingerprinting attempt to prove the ownership upon IP violation, but they are often too late to stop catastrophic damage of IP abuse and too feeble against strong adversaries. In this paper, we propose IDEA, an Inverse Domain Expert Adaptation based proactive DNN IP protection method featuring active authorization and source traceability. IDEA generalizes active authorization as an inverse problem of domain adaptation. The multi-adaptive optimization is solved by a mixture-of-experts model with one real and two fake experts. The real expert re-optimizes the source model to correctly classify test images with a unique model user key steganographically embedded. The fake experts are trained to output random prediction on test images without or with incorrect user key embedded by minimizing their mutual information (MI) with the real expert. The MoE model is knowledge distilled into a unified protected model to avoid leaking the expert model features by maximizing their MI with additional multi-layer attention and contrastive representation loss optimization. IDEA not only prevents unauthorized users without the valid key to access the functional model, but also enable the model owner to validate the deployed model and trace the source of IP infringement. We extensively evaluate IDEA on five datasets and four DNN models to demonstrate its effectiveness in authorization control, culprit tracing success rate, and robustness against various attacks.
@@ I Introduction
In view of the rampant DNN intellectual property (IP) theft, various protection methods have been proposed in recent years. The approaches can be categorized into passive and active protection methods. Passive protection methods, such as watermarking [6, 7, 8, 9, 10, 11, 12] and fingerprinting [13, 14, 15, 16], aim at detecting the ownership of the deployed DNN upon IP infringement, whereas active protection methods, such as [17, 18, 19, 20, 21], aim at preventing IP violation proactively. The latter limits unauthorized users without a valid key from accessing the full utility of the deployed model by multi-task optimization. Constrained by multiple optimization objectives, these optimization techniques mainly alter the decision boundaries of the last and/or penultimate layers. As the feature distribution in the shallow layers remains almost identical for both authorized and unauthorized samples, they can be easily exploited for successful fine-tuning attacks.
@@
In this paper, we propose IDEA – an active DNN IP protection method based on Inverse Domain Expert Adaptation. It enables the owner to proactively lock the model’s functionality from access by unauthorized users without a valid user key. Even if the model has been misappropriated, the owner can still claim the ownership and trace the culprit responsible for the model leakage. Our novel contributions are as follows. Unlike existing methods, IDEA treats the active authorization IP protection as an inverse problem of domain adaptation and solves it by training a mixture of experts (MoE) to minimize the mutual information (MI) between the authorized source and unauthorized target domains. A real expert is fine-tuned from the unprotected model to produce accurate prediction on authorized images. Authorized images can be produced by hiding a user key using a trained invertible Steganographic Generative Adversarial Network (SteganoGAN). The encoder of the SteganoGAN is distributed to legitimate users while the decoder is kept private by the model owner for user tracking. Two fake experts are trained to intentionally degrade the prediction accuracy on unencoded clean images and images encoded with invalid key, respectively. The three independently trained experts are inseparably fused into a coherent protected model by knowledge distillation across several selected hidden layers from shallow to depth. The stealthiness, effectiveness, fidelity and uniqueness of IDEA are evaluated by extensive experiments across five popular datasets and four network architectures. Its superiority over existing active protection methods and robustness against a wide range of attacks have also been validated.
@@
The rest of the paper is organized as follows. Related works are reviewed in Section II. Our threat model and problem formulation, background knowledge on SteganoGAN and domain expected risk bound are presented in Section III. The proposed IDEA is elaborated in Section IV. Section V presents the experimental results and comparison, followed by the latent representations visualization and security analysis in Sections VI and VII, respectively. The paper is concluded in Section VIII.
@@ II-B Active Protection
HPNN [17] embeds the encoded model into a trustworthy hardware device. In [18], the model is trained on pixel-shuffled samples to attain optimal inference performance exclusively on test samples that have been shuffled with the same secret key. ChaoWs [20] rearranges the neuron positions to encrypt a trained model, which can only be decrypted with the secret key provided by the owner. M-LOCK [23] and SSAT [21] trains the model with correctly labeled trigger-embedded training samples and wrongly labeled clean training samples so that the trained model can only accurately classify images with the pre-defined backdoor trigger. DSN [24] pastes a binary pattern trigger on training images, and utilizes a gradient reversal layer [25] to train the protected model to behave differently on images with and without the trigger. NTL [26] restricts the model’s generalization ability to a certain domain to realize active authorization.
DeepIPR [19] is the first passport-based protection scheme for access control and ownership verification. This method replaces selected normalization layers of the model with passport layers to degrade its performance. Authorized users can restore the model’s original performance by submitting a valid passport image along with the test images. PAN [27] jointly trains the target model with a passport-aware branch to avoid model structure modification; however, this method lacks the usage control in the distributed passport-free branch. TdN [28] enables irreversible ownership verification against ambiguity attacks with oracle passports. Recently, Steganographic passport [29] hides the user’s identity images into the passport with an invertible stenographic network. This approach offers robust ownership proof without necessitating extra model retraining for each new user.
@@ III-A Threat Model
Under the active protection paradigm, the defender is assumed to have complete knowledge and control over all components and training pipelines of a model before it is distributed and deployed. As the adversary can be a legal buyer who illegally redeploys the purchased model for profit beyond the contracted term, the protected model’s parameters and architecture are assumed to be transparent to the adversaries. During the verification stage, the defender can only query the suspected model with a few reserved test samples. The outputs can be analyzed to determine if it is a counterfeit copy, and trace the source of counterfeit if it is.
@@ III-B Problem Formulation
We formulate the active authorization as an inverse problem of domain adaptation. Our goal is to maintain a high inference accuracy of the protected model on the authorized (source) domain, where the input data are embedded with the correct key, while weakening its generalizability on unauthorized (target) domains with missing or incorrect keys. Given a pre-trained encoder $\mathcal{E}(\cdot,\cdot)$ and a binary key $k_{j}\in\mathcal{K}$, the following three domains of input samples can be defined.
$\vspace{-1mm}\mathcal{B}$ =$=\left\{(x,y)\mid x\sim\mathcal{P}_{X}^{\mathcal{B}},y\sim\mathcal{P}_{Y}^{\mathcal{B}}\right\},$ (1)
$\mathcal{A}_{j}$ =$=\left\{(x^{+}_{j},y)\mid x^{+}_{j}:=\mathcal{E}(x,k_{j}),x^{+}_{j}\sim\mathcal{P}_{X}^{\mathcal{A}_{j}},y\sim\mathcal{P}_{Y}^{\mathcal{A}_{j}}\right\},$ (2)
$\mathcal{N}_{j}$ =$=\left\{(x^{-}_{j},y)\mid x^{-}_{j}:=\mathcal{E}(x,k^{\ast}),x^{-}_{j}\sim\mathcal{P}_{X}^{\mathcal{N}_{j}},y\sim\mathcal{P}_{Y}^{\mathcal{N}_{j}}\right\}.\vspace{-1mm}$ (3)
@@
The benign domain $\mathcal{B}$ consists of the original samples, where $\mathcal{P}_{X}^{\mathcal{B}}$ and $\mathcal{P}_{Y}^{\mathcal{B}}$ denote the distributions of benign images and labels, respectively. The authorized domain $\mathcal{A}_{j}$ contains authorized samples that are all embedded with the correct key $k_{j}$. The noise domain $\mathcal{N}_{j}$ contains samples that are each embedded with a different random wrong key $k^{\ast}\neq k_{j}$.
@@
We consider a $K$-class image classifier $f_{\theta}(\cdot)$ trained on the training dataset $D_{\text{tr}}=\{(x_{i},y_{i})\}_{i=1}^{N_{\text{tr}}}$, with $\theta$ being the trained parameters. Instead of directly releasing the unprotected model $f_{\theta}(\cdot)$ to $M$ paying customers, the model owner further generates $M$ different protected models $\{f_{\theta_{1}}(\cdot),f_{\theta_{2}}(\cdot),\cdots,f_{\theta_{M}}(\cdot)\}$ from $f_{\theta}(\cdot)$ by embedding into each instance one of the $M$ unique user-specific keys $\mathcal{K}=\{k_{1},k_{2},\cdots,k_{M}\}$. Each protected model has special prediction behaviors on input samples from different domains.
@@
Definition 1 (Key-based Active Authorization). Given the benign domain $\mathcal{B}$, a specific binary key $k_{j}$, and the corresponding authorized and noise domains $\mathcal{A}_{j}$ and $\mathcal{N}_{j}$, respectively. A model $f_{\theta_{j}}(\cdot)$ protected by the key-based active authorization is a $k_{j}$ encoded instance derived from an unprotected model $f_{\theta}(\cdot)$ to perform well only on $\mathcal{A}_{j}$ but predicts randomized outputs on $\mathcal{B}$ and $\mathcal{N}_{j}$ like a untrained model. The protected model fulfills the following three properties:
$\mathbb{E}_{(x,y)\sim\mathcal{B}}\left[\mathbb{I}\{f_{\theta}(x)=y\}\right]-\mathbb{E}_{(x^{+}_{j},y)\sim\mathcal{A}_{j}}\left[\mathbb{I}\{f_{\theta_{j}}(x^{+}_{j})=y\}\right]<\epsilon_{1},$ (4)
$\mathbb{E}_{(x,y)\sim\mathcal{B}}\left[\mathbb{I}\{f_{\theta}(x)=y\}\right]-\mathbb{E}_{(x,y)\sim\mathcal{B}}\left[\mathbb{I}\{f_{\theta_{j}}(x)=y\}\right]>\epsilon_{2},$ (5)
$\mathbb{E}_{(x,y)\sim\mathcal{B}}\left[\mathbb{I}\{f_{\theta}(x)=y\}\right]-\mathbb{E}_{(x^{-}_{j},y)\sim\mathcal{N}_{j}}\left[\mathbb{I}\{f_{\theta_{j}}(x^{-}_{j})=y\}\right]>\epsilon_{2},$ (6)
@@
where $\mathbb{E}_{(x,y)\sim\mathcal{B}}\left[\mathbb{I}(f_{\theta}(x)=y)\right]$ represents the accuracy of the unprotected model evaluated on the benign domain. $\epsilon_{1}$ and $\epsilon_{2}$ denote two positive fractions that set the accuracy preservation and degradation thresholds, respectively.
@@
Eq. (4) specifies the fidelity requirement on authorized access, which requires the protected model $f_{\theta_{j}}(\cdot)$ to guarantee performance of no more than a trivial accuracy drop from $f_{\theta}(\cdot)$ on images encoded with the correct key $k_{j}$. Eq. (5) stipulates the effectiveness against unauthorized access, which requires the accuracy of $f_{\theta_{j}}(\cdot)$ to fall below that of the original model by a substantial margin in the absence of a valid key $k_{j}$. Eq. (6) specifies the uniqueness criterion, which calls for a unique key to affiliate with each encoded model instance. Mismatched model-key pairs will result in a substantial accuracy degradation from its source model.
Each model-key pair (i.e., $\{f_{\theta_{j}}(\cdot),k_{j}\}$) along with the unified encoder $\mathcal{E}(\cdot,\cdot)$ are distributed to the corresponding legal user. Given only black-box access to a suspected model ${f}_{\theta^{\prime}}(\cdot)$, the following definition supports ownership verification and dishonest user identification.
Definition 2 (Ownership Verification). Let $D_{q}$ denotes a benign query set sampled from the benign domain, the ownership of the model owner over the suspected model $f_{\theta^{\prime}}(\cdot)$ is verified if $\exists!k_{j}\in\mathcal{K}$ fulfills:
$\mathbb{E}_{(x,y)\sim D_{q}}\left|\mathbb{I}\{f_{\theta}(x)=y\}-\mathbb{I}\{f_{\theta^{\prime}}(\mathcal{E}(x,k_{j}))=y\}\right|<\epsilon_{1},$ (7)
$\mathbb{E}_{(x,y)\sim D_{q}}\left|\mathbb{I}\{f_{\theta^{\prime}}(x)=y\}-\mathbb{I}\{f_{\theta^{\prime}}(\mathcal{E}(x,k_{j}))=y\}\right|>\epsilon_{2}.$ (8)
Eq. (7) proves that the full functionality of the suspected model can be activated by a recorded key, and Eq. (8) rules out its possibility as an innocent model. The user who is assigned the traced key $k_{j}$ can be determined from the illegally resold or redistributed model.

### v2

@@ Abstract
Illegitimate reproduction, distribution and derivation of Deep Neural Network (DNN) models can inflict economic loss, reputation damage and even privacy infringement. Passive DNN intellectual property (IP) protection methods such as watermarking and fingerprinting attempt to prove the ownership upon IP violation, but they are often too late to stop catastrophic damage of IP abuse and too feeble against strong adversaries. In this paper, we propose IDEA, an Inverse Domain Expert Adaptation based proactive DNN IP protection method featuring active authorization and source traceability. IDEA generalizes active authorization as an inverse problem of domain adaptation. The multi-adaptive optimization is solved by a mixture-of-experts model with one real and two fake experts. The real expert is re-optimized from the source model to correctly classify authorized images embedded with the correct user key. The two fake experts are separately trained to output random prediction on test images without or with incorrect user key embedded, respectively. The constructed MoE model is knowledge distilled into a unified protected model to avoid functionality leakage. IDEA not only prevents unauthorized users without the valid key to access the functional model, but also enable secure licensee verification and reliable ownership verification. We extensively evaluate IDEA on four datasets and four DNN models to demonstrate its stealth, effectiveness, fidelity, uniqueness, and robustness against various attacks.
@@ I Introduction
In view of the rampant DNN intellectual property (IP) theft, various protection methods have been proposed in recent years. The approaches can be categorized into passive and active protection methods. Passive protection methods, such as watermarking [6, 7, 8, 9, 10, 11, 12] and fingerprinting [13, 14, 15, 16], aim at detecting the ownership of the deployed DNN upon IP infringement, whereas active protection methods [17, 18, 19, 20, 21, 22, 23] aim at preventing IP violation proactively. This work mainly focuses on active protection which safeguards the target model by locking its functionality to unauthorized usage. That is, the protected model performs normally without substantially degraded inference performance only when the correct key is present.
Research on active protection methods are in the nascent state of development. Existing active protection solutions face notable challenges, including dependence on hardware [17, 22], low fidelity [18], and limited robustness [21, 18, 19, 24]. Notably, some approaches [21, 23, 25, 26] achieve active protection through multi-task optimization, aiming for the model to perform well only on authorized samples. However, these techniques mainly alter the decision boundaries of the last and/or penultimate layers. Consequently, the feature distribution in the shallow layers remains almost identical for both authorized and unauthorized samples, making the protected model susceptible to fine-tuning attacks.
@@
In this paper, we propose IDEA – an Inverse Domain Expert Adaptation based multi-user active DNN IP protection framework. Unlike existing methods, IDEA treats the active authorization IP protection as an inverse problem of domain adaptation, and solves it by training a mixture of experts (MoE) to minimize the mutual information (MI) between authorized source and unauthorized target domains. Encoded images are produced by embedding a binary string (referred to as key) using a pretrained SteganoGAN [27]. The encoder of the SteganoGAN is distributed to all legitimate users, while the decoder is kept private by the model owner. To construct the MoE model, a real expert is fine-tuned from the unprotected model to produce accurate prediction on authorized images (encoded with the correct key), and two fake experts are trained to intentionally degrade the prediction accuracy on clean images (unencoded) and noise images (encoded with an invalid key), respectively. The three independently trained experts are inseparably fused into a coherent protected model by knowledge distillation across several selected hidden layers from shallow to deep. Consequently, only the authorized user who holds the correct key can generate authorized images that the protected model can process normally.
Our contributions are as follows.
  • We systematically define the requirements for multi-user active DNN IP protection, and propose IDEA which constructs a MoE model with one real expert and two fake experts, and further distill it into a coherent protected model.
  • We also offers two verification mechanisms: 1) Licensee verification. Authorized users can securely prove their legitimacy by simply providing encoded test images without having to divulge their secret keys; 2) Ownership verification. Even if the model has been misappropriated, the owner can still conduct black-box queries on the suspected model to claim ownership and trace the culprit responsible for the model leakage.
  • We conduct extensive evaluations across four datasets and four model architectures to demonstrates the superior performance of IDEA in terms of stealth, effectiveness, fidelity, and uniqueness. Its superiority over existing active protection methods and robustness against a wide range of attacks have also been validated.
@@
The rest of the paper is organized as follows. Related works are reviewed in Section II. Our threat model and problem formulation, background knowledge on SteganoGAN and domain expected risk bound are presented in Section III. The proposed IDEA is elaborated in Section IV. Section V presents the experimental results and comparison, followed by the latent representations visualization and security analysis in Sections VI and VII, respectively. Limitations of IDEA are discussed in Section VIII. The paper is concluded in Section IX.
@@ II-A Passive Protection
Apart from image classification tasks, passive protection methods for generative models such as generative adversarial networks (GANs) [29, 30, 31] and large language models (LLMs) [32, 33], have also been explored in recent years.
@@ II-B Active Protection
HPNN [17] embeds the encoded model into a trustworthy hardware device. NNSplitter [22] utilizes a reinforcement learning algorithm to identify a few important filters and obfuscates them to lock the target model’s normal functionality. This method relies on trusted execution environment to secure the model inference, which is not flexible and scalable without inference time overhead. In [18], the model is trained on pixel-shuffled samples to attain optimal inference performance exclusively on test samples that have been shuffled with the same secret key. This method exhibits low fidelity as its pixel shuffling operation distorts the semantic integrity of test images.
M-LOCK [25], SSAT [21], and SecureNet [23] train the model with correctly labeled trigger-embedded training samples and wrongly labeled clean training samples, such that the protected model can only accurately classify images with the pre-defined backdoor trigger. Similarly, DSN [26] pastes a binary pattern trigger on training images, and utilizes a gradient reversal layer [34] to train the protected model to behave differently on images with and without the trigger. NTL [35] restricts the model’s generalization ability to a certain domain to realize active authorization. However, these methods focus on only modifying the decision boundaries of the last and/or penultimate layers, while the feature representations in the shallow layers still leaks the functionality of the protected model.
DeepIPR [19], initially designed for DNN watermarking, can also be employed for active protection. By replacing selected normalization layers of the DNN with passport layers, the protected model performs normally only when the valid passport is presented together with the input. Following DeepIPR, several advanced passport-based protection methods [24, 36, 37] are further proposed with improved fidelity, robustness, and flexibility. Unfortunately, the passport layers can be easily located and attacked. Chen et al. [38] proposed an effective ambiguity attack against passport-based protection methods by substituting the located passport layers.
@@ III-A Threat Model
The model owner who invests significant effort and resources into training a high-performance DNN model can merchandize its model either as an online service or a distributed IP. The online revenue model allows multiple users to register and pay for inference services either on subscription or per-use basis through API calls to the remotely deployed model. It protects the model owner from the risk of IP theft and model misappropriation, but users may have concern about exposing sensitive or private information when directly submitting test samples to the model owner for predictions. Another revenue model is to sell distributed instances of a DNN model to users at a loyalty fee or blanket IP release price for conditional (e.g., non-exclusive license, non-redistribution, lawful use, etc.) unlimited use on users’ local devices for offline predictions. This IP distribution revenue model poses a severe threat to the model IP due to difficulty in enforcement of certain sale agreements. The model’s architecture and weights become fully accessible to authorized users. A malicious attacker could steal the deployed model from an authorized user and use it illegally without paying or abuse it for AI scams and frauds. Even more concerning is the possibility of a malicious authorized user who, after purchasing the model once, illegally deploys it on the cloud to provide API-based predictions for profit. Such misuse results in substantial financial losses for the owner.
Owner’s knowledge and capabilities: The owner is assumed to have full knowledge and control over all components and training pipelines of the model prior to its distribution and deployment. This enables the owner to modify the training dataset and optimization process as needed to generate multiple protected versions of the model.
Owner’s goals: Each protected model should function normally only when the correct key is provided. This ensures that authorized users who have purchased the model can use it without performance degradation, while attackers without a key or other authorized users with mismatched keys cannot use it to obtain accurate inference. Authorized users should be able to prove their right of use by presenting a few, or even a single, encoded image to the owner without revealing the secret key. In the event of illegal redeployment, the owner must be able to assert ownership of the suspected model and identify the authorized user responsible for the model’s leakage or misuse. To counteract potential modifications by attackers attempting to evade tracing, the protection method must be robust against various attacks, such as fine-tuning, model pruning, transfer learning, and reverse engineering.
@@ III-B Problem Formulation
We formulate the active authorization as an inverse problem of domain adaptation. Our goal is to maintain a high inference accuracy of the protected model on the authorized (source) domain, where the input data are embedded with the correct key, while weakening its generalizability on unauthorized (target) domains with missing or incorrect keys. Given a pretrained encoder $\mathcal{E}(\cdot,\cdot)$ and a binary key $k$ which encodes the identity information of both the owner and a specific user (e.g., “Alice-Bob”), the following three domains of input samples can be defined.
$\mathcal{B}$ =$=\left\{(x,y)\mid x\sim\mathcal{P}_{X}^{\mathcal{B}},y\sim\mathcal{P}_{Y}^{\mathcal{B}}\right\},$ (1)
$\mathcal{A}$ =$=\left\{(x^{+},y)\mid x^{+}:=\mathcal{E}(x,k),x^{+}\sim\mathcal{P}_{X}^{\mathcal{A}},y\sim\mathcal{P}_{Y}^{\mathcal{A}}\right\},$ (2)
$\mathcal{N}$ =$=\left\{(x^{-},y)\mid x^{-}:=\mathcal{E}(x,k^{\ast}),x^{-}\sim\mathcal{P}_{X}^{\mathcal{N}},y\sim\mathcal{P}_{Y}^{\mathcal{N}}\right\}.$ (3)
@@
The benign domain $\mathcal{B}$ consists of the original samples, where $\mathcal{P}_{X}^{\mathcal{B}}$ and $\mathcal{P}_{Y}^{\mathcal{B}}$ denote the distributions of benign images and labels, respectively. The authorized domain $\mathcal{A}$ contains authorized samples that are all embedded with the correct key $k$. The noise domain [^1] It should be emphasized that a noise image in IDEA context refers to an unauthorized test image with a wrong embedded key. It should not be misinterpreted as Gaussian noise or visible perturbations as the noise image may not necessarily look distorted or noisy since the wrong key is stegnographically embedded. $\mathcal{N}$ contains samples that are each embedded with a different random wrong key $k^{\ast}\neq k$.
@@
We consider an image classifier $f$ trained on the benign training dataset $D_{\text{tr}}=\{(x_{i},y_{i})\}_{i=1}^{N_{\text{tr}}}$. Instead of directly releasing the unprotected model $f$ to $M$ paying customers, the model owner further generates $M$ different protected models $\{f^{\prime}_{1},f^{\prime}_{2},\cdots,f^{\prime}_{M}\}$ from $f$ by embedding into each instance one of the $M$ unique user-specific keys $\mathcal{K}=\{k_{1},k_{2},\cdots,k_{M}\}$. Each protected model has special prediction behaviors on input samples from different domains. For brevity, the user index is omitted, and only a single protected model and corresponding user key are presented below.
@@
Definition 1 (Key-based Active Authorization). Given the benign domain $\mathcal{B}$, a specific binary key $k$, and the corresponding authorized and noise domains $\mathcal{A}$ and $\mathcal{N}$, respectively. A protected model $f^{\prime}$ is an encoded instance which performs well only on $\mathcal{A}$ but predicts randomized outputs on $\mathcal{B}$ and $\mathcal{N}$ like an untrained model. The protected model fulfills the following three properties:
$\left|\operatorname{Acc}\{f,\mathcal{B}\}-\operatorname{Acc}\{f^{\prime},\mathcal{A}\}\right|<\epsilon_{1},$ (4)
$\operatorname{Acc}\{f,\mathcal{B}\}-\operatorname{Acc}\{f^{\prime},\mathcal{B}\}>\epsilon_{2},$ (5)
$\operatorname{Acc}\{f,\mathcal{B}\}-\operatorname{Acc}\{f^{\prime},\mathcal{N}\}>\epsilon_{2},$ (6)
@@
where $\operatorname{Acc}\{\cdot,\cdot\}$ evaluates the provided model’s performance on a specific domain or test dataset. $\epsilon_{1}$ and $\epsilon_{2}$ denote two positive fractions that set the accuracy preservation and degradation thresholds, respectively.
@@
Eq. (4) specifies the fidelity requirement on authorized access, which requires the protected model $f^{\prime}$ to guarantee performance on images encoded with the correct key $k$. Eq. (5) stipulates the effectiveness against unauthorized access, which requires the accuracy of $f^{\prime}$ to fall below that of the original model by a substantial margin in the absence of $k$. Eq. (6) specifies the uniqueness criterion, which calls for a unique key to affiliate with each encoded model instance. Mismatched model-key pairs will result in a substantial accuracy degradation from its source model. Each model-key pair (i.e., $\{f^{\prime},k\}$) along with the unified encoder $\mathcal{E}$ are distributed to the corresponding legal user.